OpenStudy (markchernioglo):

Please help me with these Quadratic Equations questions, I have an exam tomorrow, and I want to pass with an A+! Thanks! :)

OpenStudy (markchernioglo):

Which is a quadratic model for the given data? x: 0.9 2.0 3.2 4.1 y: 2 8 18 32 A. y=2.2x^2-1.8x+2.1 B. y= 2.1x^2+2.2x-1.8 C. y= -1.8x^2+2.1x+2.2

OpenStudy (princeharryyy):

easy method-plug in the values of x and find y using equations. hard method- take y = ax^2+bx+c as the gen equation. plug in values three times find a b c

OpenStudy (3mar):

and the easiest way,@markchernioglo is to do such that: https://www.desmos.com/calculator/qxtkdg0kw5

OpenStudy (markchernioglo):

Thank you both prince and 3mar! :)

OpenStudy (3mar):

Don't mention it! Any Help... Any Time... What matters is that you got the idea!